OpenStudy (anonymous):

Five years ago a girl was one-half as old as her brother was then, and now she is three years younger than he is. What are their present ages? girl is 7; brother is 10 girl is 8; brother is 11 girl is 9; brother is 12 not sure but i think its a

OpenStudy (anonymous):

girl is 8; brother is 11

OpenStudy (anonymous):

just subtract 5 from each number 8-5 = 3, 11-5 = 6. 3/6 = 1/2

OpenStudy (anonymous):

Refer to the Mathematica attachment.